OpenStudy (anonymous):

How do you solve this using Gaussian elimination? x + 2y + 3z = - 6 2x - 3y - 4z = 15 3x + 4y + 5z = - 8

OpenStudy (anonymous):

I first took the first equation and multiplied it by -2 to get - 2x -4y -6z = 12 -2x -4y -6z = 12 2x - 3y -4z = 15 Those added to -7y -10z = 27